Calculate the heat required to melt 7.87 g of benzene at its normal melting point.

Heat of fusion (benzene) = 9.92 kJ/mol
1
Expert's answer
2017-12-08T07:02:17-0500
Mollar mass (C6H6) = 78 g/mol
n = 7.87 / 78 = 0.101 (mol)
Q = 0.101 * 9.92 = 1.002 (kJ)